# Beyond MOT: Semantic Multi-Object Tracking

### v1

@@ Abstract
Current multi-object tracking (MOT) aims to predict trajectories of targets (i.e., “where”) in videos. Yet, knowing merely “where” is insufficient in many crucial applications. In comparison, semantic understanding such as fine-grained behaviors, interactions, and overall summarized captions (i.e., “what”) from videos, associated with “where”, is highly-desired for comprehensive video analysis. Thus motivated, we introduce Semantic Multi-Object Tracking (SMOT), that aims to estimate object trajectories and meanwhile understand semantic details of associated trajectories including instance captions, instance interactions, and overall video captions, integrating “where” and “what” for tracking. In order to foster the exploration of SMOT, we propose BenSMOT, a large-scale Benchmark for Semantic MOT. Specifically, BenSMOT comprises 3,292 videos with 151K frames, covering various scenarios for semantic tracking of humans. BenSMOT provides annotations for the trajectories of targets, along with associated instance captions in natural language, instance interactions, and overall caption for each video sequence. To our best knowledge, BenSMOT is the first publicly available benchmark for SMOT. Besides, to encourage future research, we present a novel tracker named SMOTer, which is specially designed and end-to-end trained for SMOT, showing promising performance. By releasing BenSMOT, we expect to go beyond conventional MOT by predicting “where” and “what” for SMOT, opening up a new direction in tracking for video understanding. Our BenSMOT and SMOTer will be released.
@@ 1 Introduction
Thus motivated, in this work we introduce a new type of MOT task, dubbed Semantic Multi-Object Tracking (SMOT) [^1] Here by “semantic”, we emphasize high-level trajectory-based activity understanding in videos in the context of tracking, instead of category as in semantic segmentation., which aims to expand scope of MOT beyond merely predicting trajectories to capture rich semantic details of objects in videos. Specifically, besides trajectory estimation, SMOT comprises three additional trajectory-associated semantic understanding tasks, including instance captioning, instance interaction recognition, and video captioning, as illustrated in Fig. 1. In particular, instance captioning aims at describing objects and their behaviors in human language, answering “what are the objects doing”; instance interaction recognition is to capture relations between objects, answering “what are the relations between objects”; video captioning provides the overall scenario understanding based on trajectories, answering “what is happening”. Note that, all these three additional tasks are associated with object trajectories in videos. Ultimately, the goal of SMOT is to integrate “where” (i.e., object trajectories) and “what” (i.e., trajectory-associated semantics) for video understanding, going beyond MOT predicting only trajectories (see again Fig. 1). Compared to conventional MOT, our SMOT is a multimodal task across vision and language, which is more challenging yet practical. It is worthy to notice that, SMOT is a natural extension of MOT from vision to vision-language. Despite sharing some similarities with video captioning [23], SMOT differs in that it aims at language understanding of dense target trajectories from videos in the context of instance-level tracking.
@@ 3.3 Data Annotation
To ensure the high-quality annotations, we adopt a multi-round mechanism. Specifically, each video will be first manually annotated by a few volunteers who are familiar with the task and an expert working on related problems. After this initial round, all annotations, including bounding box, instance caption, interaction, and video caption, of the video will be sent to a validation team former by more than two experts for inspection. If initial annotations are not unanimously agreed by the experts, they will be returned to the original labeling team for refinement using the feedback from the validation team. We repeat this procedure for each video until the annotations of all video sequences in BenSMOT are completed. An annotation example in our BenSMOT is shown in Fig. 1 (b), and more can be seen in supplementary material.
(a) (a) Trajectory length distribution
(b) (b) Sequence length distribution
(c) (c) Interactions count distribution
(d) (d) Inst caption length distribution
(e) (e) Video caption length distribution
@@ 3.4 Dataset Split and Evaluation Metric
Evaluation Metric. We apply multiple metrics on BenSMOT for evaluation. Specifically, to assess the performance in object trajectory estimation, we employ higher order tracking accuracy (HOTA), association accuracy (AssA), detection accuracy (DetA), and localization accuracy (LocA) by following [30], CLEAR metrics [3] including multiple object tracking accuracy (MOTA), false positives (FP), false negatives (FN), and ID switches (IDs), and ID metrics [35] containing identification precision (IDP), identification recall (IDR) and related F1 score (IDF1). For instance and video captioning, we follow existing methods (e.g.[27, 37]) and employ BLEU [34], ROUGE [26], METEOR [2], and CIDEr [42] for thorough evaluation. Lastly, to assess performance of interaction recognition, we leverage classic metrics such as Precision (Prcn), Recall (Rcll), and F1, which are widely used in the context of classification tasks. For a more comprehensive understanding of these metrics, please refer to [30, 3, 35, 34, 26, 2, 42].
@@ 4.2 Feature Fusion
Given image and target features from last stage in Sec. 4.1, we design two feature fusion modules, including attention-based Video Fusion Module (VFM) and concatenation-based Trajectory Fusion Module (TFM). VFM aims to merge image-level features into overall video features while TFM focuses on integrating target features from each frame into target trajectory features.
@@ 4.3 Semantic Understanding
where ${\tt Dec}_{\text{Vid}}(\cdot)$ is a video text decoder for video captioning (please kindly refer to supplementary material for details), and ${\tt Linear}(\cdot)$ a linear projection.
@@
$R_{\text{Int}}^{j,k}={\tt MLP}({\tt CA}(\tilde{T_{j}},\tilde{T_{k}}),$ (5)
@@ 5 Experiments
Method HOTAnormal-$\uparrow$ AssAnormal-$\uparrow$ DetAnormal-$\uparrow$ LocAnormal-$\uparrow$ MOTAnormal-$\uparrow$ FNnormal-$\downarrow$ FPnormal-$\downarrow$ IDsnormal-$\downarrow$ IDRnormal-$\uparrow$ IDPnormal-$\uparrow$ IDF1normal-$\uparrow$
@@
Method BLEUnormal-$\uparrow$ ROUGEnormal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ BLEUnormal-$\uparrow$ ROUGEnormal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ Precnormal-$\uparrow$ Rcllnormal-$\uparrow$ F1normal-$\uparrow$
@@
BLEUnormal-$\uparrow$ ROUGEnormal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ BLEUnormal-$\uparrow$ ROUGEnormal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ Prcnnormal-$\uparrow$ Rcllnormal-$\uparrow$ F1normal-$\uparrow$
@@ 5.2 Comparison on Semantic Understanding
Due to limited space, more results can be seen in supplementary material.
(a) (a) Comparison results on object tracking
(b) (b) Comparison results on video captioning
(c) (c) Comparison results on instance captioning
@@ 5.3 Ablation Study
Impact of different feature fusion. Feature fusion is a crucial component in our SMOTer for generating overall video and enhanced trajectory features. In this work, we study four types of feature fusion strategies for VFM and TFM, including attention-based fusion (using cross-attention), MLP-based fusion (using MLP module), concatenation-based fusion, and addition-based fusion. Due to space limitation, please refer to supplementary material for their detailed architectures. To assess the impact for VFM, we compare the performance on video caption generation because the fused video feature from VFM is used for this sub-task, while for TFM, we measure the results on instance captioning and interaction recognition since these two sub-tasks rely the fused trajectory features. Tab. 4 shows the results of different fusion mechanisms for VFM and TFM. From Tab. 4, we can see that attention-based fusion works generally better for VFM by achieving best ROUGE score of 0.261 and METEOR score of 0.223 and second best BLEU score of 0.245 and CIDEr score of 0.343 for video captioning. For TFM, concatenation-based fusion shows more superior results by obtaining best scores on ROUGE, METEOR, CIDEr for instance captioning and on Precision and F1 for interaction recognition.
@@
$\tau_{p}$ HOTAnormal-$\uparrow$ MOTAnormal-$\uparrow$ IDF1normal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ METEORnormal-$\uparrow$ CIDErnormal-$\uparrow$ Prcnnormal-$\uparrow$ Rcllnormal-$\uparrow$ F1normal-$\uparrow$
@@ 5.4 Discussion
Semantic understanding for improving tracking. SMOT aims to go beyond just tracking by incorporating additional semantic understanding tasks including trajectory-associated instance/video captioning and interaction recognition. We observe that these extra goals greatly improve the tracking performance. For example, as shown in Tab. 2, we can see that joint training of different tasks leads to better performance of SMOTer (71.98%/77.71% in HOTA/MOTA) in tracking compared to its baseline ByteTrack (68.84%/73.87 in HOTA/MOTA). The possible reason is that SMOTer needs to comprehensively understand the video content for semantic multi-object tracking, and the the learned semantic knowledge is able to help distinguish different object trajectories.
Challenge in SMOT. One reason why SMOT is challenging lies in the complicated behaviors of targets. This complexity often results in lengthy instance captions. For instance, during a basketball game, a player may go through various states such as dribbling, shooting, running, and defending. However, models often overlook certain states during the understanding process. In BenSMOT, the average video length is approximately 23 seconds, with objects typically persisting throughout, leading to an average trajectory length of around 21 seconds. As a result, the average instance caption length exceeds 35 words. This brings a significant challenge for SMOT and can be a direction to improve in future.

### v4

@@ Abstract
Current multi-object tracking (MOT) aims to predict trajectories of targets (i.e., “where”) in videos. Yet, knowing merely “where” is insufficient in many crucial applications. In comparison, semantic understanding such as fine-grained behaviors, interactions, and overall summarized captions (i.e., “what”) from videos, associated with “where”, is highly-desired for comprehensive video analysis. Thus motivated, we introduce Semantic Multi-Object Tracking (SMOT), that aims to estimate object trajectories and meanwhile understand semantic details of associated trajectories including instance captions, instance interactions, and overall video captions, integrating “where” and “what” for tracking. In order to foster the exploration of SMOT, we propose BenSMOT, a large-scale Benchmark for Semantic MOT. Specifically, BenSMOT comprises 3,292 videos with 151K frames, covering various scenarios for semantic tracking of humans. BenSMOT provides annotations for the trajectories of targets, along with associated instance captions in natural language, instance interactions, and overall caption for each video sequence. To our best knowledge, BenSMOT is the first publicly available benchmark for SMOT. Besides, to encourage future research, we present a novel tracker named SMOTer, which is specially designed and end-to-end trained for SMOT, showing promising performance. By releasing BenSMOT, we expect to go beyond conventional MOT by predicting “where” and “what” for SMOT, opening up a new direction in tracking for video understanding. We will release BenSMOT and SMOTer at https://github.com/Nathan-Li123/SMOTer.
@@ 1 Introduction
Thus motivated, in this work we introduce a new type of MOT task, dubbed Semantic Multi-Object Tracking (SMOT) [^1] Here by “semantic”, we emphasize high-level trajectory-based activity understanding in videos in the context of tracking, instead of category as in semantic segmentation., which aims to expand scope of MOT beyond merely predicting trajectories to capture rich semantic details of objects in videos. Specifically, besides trajectory estimation, SMOT comprises three additional trajectory-associated semantic understanding tasks, including instance captioning, instance interaction recognition, and video captioning, as illustrated in Fig. 1. In particular, instance captioning aims at describing objects and their behaviors in human language, answering “what are the objects doing”; instance interaction recognition is to capture relations between objects, answering “what are the relations between objects”; video captioning provides the overall scenario understanding based on trajectories, answering “what is happening”. Note that, all these three additional tasks are associated with object trajectories in videos. Ultimately, the goal of SMOT is to integrate “where” (i.e., object trajectories) and “what” (i.e., trajectory-associated semantics) for video understanding, going beyond MOT predicting only trajectories (see again Fig. 1). Compared to conventional MOT, our SMOT is a multimodal task across vision and language, which is more challenging yet practical. It is worthy to notice that, SMOT is a natural extension of MOT. Despite sharing some similarities with video captioning [23], SMOT differs in that it aims at language understanding of dense target trajectories from videos in the context of instance-level tracking.
@@ 3.3 Data Annotation
To ensure the high-quality annotations, we adopt a multi-round mechanism. Specifically, each video will be first manually annotated by a few volunteers who are familiar with the task and an expert working on related problems. After this initial round, all annotations of the video will be sent to a validation team former by more than two experts for inspection. If initial annotations are not unanimously agreed by the experts, they will be returned to the original labeling team for refinement using the feedback from the validation team. We repeat this procedure for each video until the annotations of all video sequences in BenSMOT are completed. An annotation example in our BenSMOT is shown in Fig. 1 (b), and more can be found in supplementary material.
@@ 3.4 Dataset Split and Evaluation Metric
Evaluation Metric. We apply multiple metrics on BenSMOT for evaluation. Specifically, to assess the performance in object trajectory estimation, we employ higher order tracking accuracy (HOTA), association accuracy (AssA), detection accuracy (DetA), and localization accuracy (LocA) by following [30], CLEAR metrics [3] including multiple object tracking accuracy (MOTA), false positives (FP), false negatives (FN), and ID switches (IDs), and ID metrics [35] containing identification precision (IDP), identification recall (IDR) and related F1 score (IDF1). For instance and video captioning, we follow existing methods (e.g.[27, 37]) and employ BLEU [34], ROUGE [26], METEOR [2], and CIDEr [42] for thorough evaluation. Lastly, to assess performance of interaction recognition, we leverage classic metrics such as Precision (Prcn), Recall (Rcll), and F1, which are widely used in the context of classification tasks.
@@ 4.2 Feature Fusion
Given image and target features from previous stage in Sec. 4.1, we design two feature fusion modules, including attention-based Video Fusion Module (VFM) and concatenation-based Trajectory Fusion Module (TFM). VFM aims to merge image-level features into overall video features while TFM focuses on integrating target features from each frame into target trajectory features.
@@ 4.3 Semantic Understanding
where ${\tt Dec}_{\text{Vid}}(\cdot)$ is a video text decoder for video captioning (please kindly refer to supplementary material for details), and ${\tt Linear}(\cdot)$ is a linear projection.
@@
$R_{\text{Int}}^{j,k}={\tt MLP}({\tt CA}(\tilde{T_{j}},\tilde{T_{k}})),$ (5)
@@ 5 Experiments
[…]
@@
[…]
@@
[…]
@@ 5.3 Ablation Study
Impact of different feature fusion. Feature fusion is a crucial component of SMOTer for generating overall video and enhanced trajectory features. In this work, we study four types of feature fusion strategies for VFM and TFM, including attention-based fusion (using cross-attention), MLP-based fusion (using MLP module), concatenation-based fusion, and addition-based fusion. Please refer to supplementary material for detailed architectures. To assess the impact for VFM, we compare the performance on video caption generation because the fused video feature from VFM is used for this sub-task, while for TFM, we measure the results on instance captioning and interaction recognition since these two sub-tasks rely on the fused trajectory features. Tab. 4 shows the results of different fusion mechanisms for VFM and TFM. We can see that attention-based fusion works generally better for VFM by achieving best ROUGE score of 0.261 and METEOR score of 0.223 and second best BLEU score of 0.245 and CIDEr score of 0.343 for video captioning. For TFM, concatenation-based fusion shows more superior results by obtaining best scores on ROUGE, METEOR, CIDEr for instance captioning and on Precision and F1 for interaction recognition.
@@
[…]
@@ 5.4 Discussion
Semantic understanding for improving tracking. SMOT aims to go beyond just tracking by incorporating additional semantic understanding tasks including trajectory-associated instance/video captioning and interaction recognition. We observe that these extra goals greatly improve the tracking performance. For example, as shown in Tab. 2, we can see that joint training of different tasks leads to better performance of SMOTer (71.98%/77.71% in HOTA/MOTA) in tracking compared to its baseline ByteTrack (68.84%/73.87 in HOTA/MOTA). The possible reason is that SMOTer needs to comprehensively understand the video content for semantic multi-object tracking, and the learned semantic knowledge helps distinguish different object trajectories.
Challenge in SMOT. One reason why SMOT is challenging lies in the complicated behaviors of targets, often leading to lengthy instance captions. For instance, a basketball player can be dribbling, shooting, and defending in a game. Models often miss some states in understanding. In BenSMOT, the average video length is about 23 seconds, with objects typically present for around 21 seconds, resulting in instance captions averaging over 35 words. This complexity poses a challenge for SMOT and indicates a potential area for future improvement.
@@ 6 Conclusion
Acknowledgements. Heng Fan was not supported by any fund for this work.